OpenStudy (anonymous):

solve by factoring. 5x^2+8x-36=0

OpenStudy (nathan917):

Simplifying 5x2 + 8x + -36 = 0 Reorder the terms: -36 + 8x + 5x2 = 0 Solving -36 + 8x + 5x2 = 0 Solving for variable 'x'. Factor a trinomial. (-18 + -5x)(2 + -1x) = 0 Subproblem 1 Set the factor '(-18 + -5x)' equal to zero and attempt to solve: Simplifying -18 + -5x = 0 Solving -18 + -5x = 0 Move all terms containing x to the left, all other terms to the right. Add '18' to each side of the equation. -18 + 18 + -5x = 0 + 18 Combine like terms: -18 + 18 = 0 0 + -5x = 0 + 18 -5x = 0 + 18 Combine like terms: 0 + 18 = 18 -5x = 18 Divide each side by '-5'. x = -3.6 Simplifying x = -3.6 Subproblem 2 Set the factor '(2 + -1x)' equal to zero and attempt to solve: Simplifying 2 + -1x = 0 Solving 2 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -1x = 0 + -2 Combine like terms: 2 + -2 = 0 0 + -1x = 0 + -2 -1x = 0 + -2 Combine like terms: 0 + -2 = -2 -1x = -2 Divide each side by '-1'. x = 2 Simplifying x = 2 Solution x = {-3.6, 2}

OpenStudy (anonymous):

For some reason this doesnt look right...

OpenStudy (anonymous):

Im attaching a screen shot.

OpenStudy (anonymous):

5x^2+8x-36=0 5x^2+18x-10x-36=0 5x(x-2)+18(x-2)=0 (5x+18)(x-2)=0 hence x=-18/5 or x=2

OpenStudy (anonymous):

@leena1996 is correct, the work is shown properly

OpenStudy (anonymous):

@marcybaby @nathan917 simpler ways must be implied to solve such factoring sums. The basic trick here is for example say ax^2+bx+c=0 is an equation, then let the products of any 2 terms be ac and the sum be b. Now think what can be the 2 numbers whose product and sum be ac and b respectively. For eg in the given question, 18,-10 are the terms where Sum b=8= 18+(-10)=8 and product ac=18*(-10)=-180=-36*5.....hope u got it.......

OpenStudy (anonymous):

thanks @galacticwavesXX

OpenStudy (nathan917):

I blame the calculator i used... Works every time..

OpenStudy (anonymous):

@nathan917 what do u mean?

OpenStudy (anonymous):

thank you @leena1996 !!! I got it!

OpenStudy (anonymous):

welcome....@marcybaby